SATURATION OF THE NO_{2} ν_{1} + ν_{3} AND THE CH_{4} ν_{3} TRANSITIONS IN HELIUM NANODROPLETS

\underline{ROBERT R. FEHNEL}, KEVIN K. LEHMANN, Department of Chemistry, University of Virginia, Charlottesville VA, 22904-4319.

We will report our findings of quantitative study of saturation of IR transition in Helium nanodroplets. The excitation source is an Argos cw PPLN OPO (Aculight) that is focused to a spot of 27 μm, producing a pear power of 87 kW/cm^{2}. We have also used a “Perry Cell” multi pass cell to study of the effect of extending the interaction time. We will present results for a study of the 2905.566 cm^{−1} R(0) transition of NO_{2} and the 3029.07 cm^{−1} R(0) transition of CH_{4}.